OpenStudy (anonymous):

can someone help me? I ask you to be very detailed, I´m just learning. limit(4x-1)/(3x+2) x=>infinit

OpenStudy (anonymous):

= 4/3

OpenStudy (anonymous):

thank you, but I just started to study this subject. I need a detailed answer.

OpenStudy (anonymous):

ok, factor 4x out from numerator, factor 3x out from denominator, ypu have \[\lim \frac{ 4x( 1 - \frac{ 1 }{ 4x }) }{ 3x ( 1+\frac{ 2 }{ x })}\]

OpenStudy (anonymous):

cancel out x you get lim 4/3 since lim 1/4x---> 0 and lim 2/x---> 0 kapeesh?

OpenStudy (anonymous):

wouldn´t the denominator be \[3x(1+2/3x)\] ?